# FAID: Fine-Grained AI-Generated Text Detection Using Multi-Task Auxiliary and Multi-Level Contrastive Learning

### v2

@@ Abstract
The growing collaboration between humans and AI models in generative tasks has introduced new challenges in distinguishing between human-written, LLM-generated, and human–LLM collaborative texts. In this work, we collect a multilingual, multi-domain, multi-generator dataset FAIDSet. We further introduce a fine-grained detection framework FAID to classify text into these three categories, and also to identify the underlying LLM family of the generator. Unlike existing binary classifiers, FAID is built to capture both authorship and model-specific characteristics. Our method combines multi-level contrastive learning with multi-task auxiliary classification to learn subtle stylistic cues. By modeling LLM families as distinct stylistic entities, we incorporate an adaptation to address distributional shifts without retraining for unseen data. Our experimental results demonstrate that FAID outperforms several baselines, particularly enhancing the generalization accuracy on unseen domains and new LLMs, thus offering a potential solution for improving transparency and accountability in AI-assisted writing.
@@ 1 Introduction
LLMs have evolved from an assistant tool to a creator or initiator, from helping polish papers and generate code to initiate proposals and draft essays, while humans increasingly serve as optimizers and reviewers. In such deeply collaborative human–LLM settings, measuring human contribution becomes challenging, yet clarifying authorship is critical for accountability and transparency, particularly in educational and academic contexts Wang et al. (2025). This work aims to identify human involvement by distinguishing the origin of a given text (in a multilingual context): (a) fully LLM-generated, (b) fully human-written, or (c) collaboratively produced.
@@
Numerous studies have explored multilingual LLM-generated text detection, but most have focused either on binary detection, i.e., human vs. LLM Wang et al. (2024d, b), or on fine-grained detection limited to English Abassy et al. (2024); Koike et al. (2024); Wang et al. (2023); Zhang et al. (2024a). Moreover, both struggle with generalization to unseen domains, languages, and LLM generators Wang et al. (2024c); Li et al. (2024).
@@
Our work aims to bridge this gap by (i) collecting a multilingual, multi-domain and multi-generator dataset, FAIDSet, for fine-grained detection, i.e., identifying a text into three categories: LLM-generated, human-written, and human–LLM collaborative; and (ii) introducing a detection framework FAID to improve generalization performance.
@@
Our dataset FAIDSet focuses on the academic field including paper abstracts, student theses, and scholarly reports, and contains generation by a variety of families of LLMs, e.g., GPT, Gemini, DeepSeek, and Llama OpenAI (2024); Team et al. (2024); DeepSeek-AI (2024); Dubey et al. (2024).
@@
Our detection framework FAID treats each LLM as a distinct author, learning specific features in the hidden space to differentiate different authors, instead of classifying based on hand-crafted stylistic features. We achieve this by optimizing a language encoder with multi-level author relationship (e.g., the stylistic similarity between texts from the same LLM family is greater than the that between a human and an LLM) to capture author-specific distinguishable signals of an input text using contrastive learning, along with the task of fine-tuning a classifier to recognize the input text’s origin. This multitask learning process forces the encoder to reorganize the hidden space, so that the representations of texts by the same authors are distributed closer than such by different authors.
@@
In practice, given that generations produced by LLMs from same company tend to have similar writing styles due to similar model architecture, training data, and training strategy (see Appendix D for more detail), we consider each LLM family as an “author.” This can also help the detector acquire prior knowledge of future LLMs from the same company. Our experiments show that FAID consistently outperforms other baseline detectors in both in-domain and out-of-domain evaluations. Our contributions can be summarized as follows:
@@
  • We collect a new multilingual, multi-domain, multi-generator dataset for fine-grained LLM-generated text detection with 84k examples.
@@
  • We propose a detection framework FAID to improve generalization performance in unseen domains and generators by capturing subtle stylistic features in the hidden representation.
@@
  • We show that FAID outperforms baseline detectors, particularly over unseen domains and generators. Meanwhile, the nature of FAID allows us to assess the stylistic proximity of a given text to other texts from our database, labeled with ground truth authorship labels.
@@ 2 Background and Related Work
Advancements in LLMs have fundamentally reshaped the process of text production and refinement. Rather than originating every word independently, people increasingly assume the roles of post-editors or reviewers, intervening after an initial draft has been generated by an LLM. This collaborative paradigm extends across diverse domains, including academic publishing, journalism, education, and social media, thus making hybrid human–AI authorship an emerging norm Cheng et al. (2024); Lee et al. (2022).
Such a shift calls question into binary authorship detection: human vs. AI. Texts can be fully human-written, fully AI-generated, or collaborative text (e.g., human-written, AI-polished; AI-generated, human-edited; and deeply mixedArtemova et al. (2025). Addressing these concerns necessitate fine-granularity authorship detection, and even tracing back to a specific LLM family, in order to assess the extent of human contribution Hutson (2025). Ensuring transparent disclosure of LLM involvement is critical for upholding research honesty and preserving the integrity of research. In response to these challenges, we collect a new dataset, FAIDSet, and propose a detection framework, FAID, which generalizes well to new domains, languages, and models, achieving consistently high accuracy and reliability.
@@ 2.1 Fine-Grained AI-generated Text Datasets
Many prior studies have explored fine-grained AI-generated text datasets across various forms of human–AI collaboration, e.g., MixSet (Zhang et al., 2024b), and Beemo (Artemova et al., 2025). See Table 3 in the Appendix, for a discussion of more datasets and detailed information regarding the label space, and the coverage of domains, languages and LLMs for each one.
However, all these studies were limited to English. A substantial gap remains in the availability of large-scale, fine-grained multilingual LLM-generated text detection datasets. To bridge this gap, we collect a multilingual fine-grained LLM-generated text detection dataset, which encompasses 84k texts generated by the latest LLMs and includes diverse forms of human–LLM collaborative generations. This dataset can facilitate the development of more robust and generalizable detection models that are capable of handling complex multilingual collaborative scenarios.
2.2 Generalization of AI-Generated Text Detection
A recent study, M4GT-Bench, Wang et al. (2024c) has highlighted a persistent challenge for both binary and fine-grained AI-generated text detection: poor generalization to unseen domains, languages, and generators. Many detection methods have shown a significant drop in performance on out-of-distribution data, underscoring the difficulty of building robust detectors to deal with real-world scenarios and evolving LLMs outputs.
Various approaches have been proposed to improve generalization. OUTFOX Koike et al. (2024) leverages adversarial in-context learning, dynamically generating challenging examples to enhance robustness, but still faces limitations in domain transferability and computational efficiency. LLM-DetectAIve Abassy et al. (2024) adopts fine-grained classification and incorporates domain-adversarial training to reduce overfitting; yet, its generalization to unseen domains and generators remains limited; moreover, its current version lacks multilingual support. SeqXGPT Wang et al. (2023) focuses on sentence-level detection by combining log-probabilities with convolutional and self-attention mechanisms, which helps to capture subtle mixed-content signals and improves generalization across input styles. However, its reliance on specific model features and limited semantic representation constrains adaptability to new generators and domains.
Finally, the DeTeCtive framework Guo et al. (2024), introduces multi-level contrastive learning to better capture writing-style diversity and enhance generalizability, especially for out-of-distribution scenarios, but it primarily focuses on binary classification and does not fully address human–AI collaborative texts.
@@ 2.3 Contrastive Learning for AI-Generated Text Detection
Contrastive learning has been extensively used to improve sentence representation by pulling semantically similar sentences closer and pushing dissimilar ones apart. SimCSE treats a sentence subjected to dropout noise as a semantically similar counterpart (i.e., a positive pair) and trains the encoder to minimize the distance between the original and the noise sentence. It further leverages natural language inference pairs, considering entailment pairs as positives and contradiction pairs as hard negatives (Gao et al., 2021). It is then trained to maximize the separation between negative pairs in the embedding space. DeCLUTR Giorgi et al. (2021) constructed positive pairs by extracting different spans from the same texts, and by sampling negative pairs from different texts.
We adopt the same core idea, but reorganize the latent space based on writing style, pulling human-written texts to cluster together while remaining distant from LLM-generated ones. Similarly to semantic textual similarity tasks, where sentence similarity ranges from 0 to 5 to reflect varying degrees of semantic overlap, we incorporate ordinal regression into our framework to model the degree of human involvement in collaborative texts, ranging from 0 (fully LLM-generated) to 1 (fully human-written).
Another work, DeTeCtive Guo et al. (2024), also took advantage of contrastive learning and was used in binary AI-generated text detection. Based on a multi-task framework, it was trained to learn writing-style diversity using a multi-level contrastive loss, and an auxiliary task of classifying the source of a given text (human vs. AI) to capture distinguishable signals. For inference, their pipeline encoded the input text as a hidden vector and applied dense retrieval to match the cluster, depending on the stylistic similarity with respect to a database of previously indexed training features. Additionally, instead of retraining the model when faced with new data, they encoded new data with the trained encoder to obtain embeddings, and then added them to the feature database to augment. This largely improved the generalizability to unseen domains and new generators.
However, this approach distinguishes only between two categories of text (human-written vs. LLM-generated) while overlooking the increasingly prevalent class of human–LLM collaborative texts. Our work bridges this gap with the aim of enhancing generalization performance for fine-grained LLM-generated text detection.
@@ 3 FAIDSet
We collected a new multilingual, multi-domain and multi-generator LLM-generated text dataset, FAIDSet. It contains texts generated by LLMs, written by humans and collaborated by both humans and LLMs, resulting in a total of 84k examples; see Appendix B for more details about FAIDSet.
FAIDSet covers two domains: student theses and paper abstracts, where identifying authorship is critical, across two languages (Vietnamese and English). We collected student theses from the database of Hanoi University of Science and Technology, and paper abstracts from arXiv and Vietnam Journals Online [^1] https://vjol.info.vn/ (VJOL).
@@
Diverse Prompt Strategies. We generated data with diverse tones and contexts while ensuring the accuracy of the content. Depending on the data source and context, we crafted prompts to create varied outputs suitable for different real-world scenarios. We generated responses with different tones using prompts such as “You are an IT student…” and “…who are very familiar with abstract writing…”. See Appendix B.2 for the full list. While FAIDSet does not capture the full natural diversity of in-the-wild LLMs outputs, the controlled setup enables us to systematically model stylistic and collaborative signals across multiple languages and families. This makes FAIDSet both a reproducible training resource and a benchmarking corpus: it provides reliable supervision for developing detectors while also serving as a testbed for assessing generalization to more diverse, unseen scenarios.
Quality Control. To avoid bad machine-generated texts, which can introduce remarkably distinguishable signals, we performed quality control by randomly sampling 10-20 instances for each domain, source, and LLM generator. Manual inspection focused on fluency, coherence and factual plausibility. Overall, the generated texts demonstrated high linguistic quality with most outputs being fluent and logically reasonable. Nevertheless, occasional issues such as repetitive phrasing, incomplete reasoning or overuse of formal expressions were observed. In that cases, we adjusted the prompts (e.g., by specifying desired length or style) or refined generation parameters to improve diversity and logical consistency. After these adjustments, the quality across generators and domains was found to be stable and satisfactory.
@@ 4 Methodology
The human–LLM collaborative category involves a range of interaction between humans and LLM systems, such as human-written, LLM-polished; human-initiated, LLM-continued; human-written, LLM-paraphrased, etc. Given the growing variety and complexity (e.g., deeply mixed text) of collaborative patterns, this is not exhaustive. Instead, we consolidate all forms of human–LLM collaboration into a single label to maintain practical simplicity and model generalizability. This reflects real-world usage, where using LLM tools to enhance clarity or expression is increasingly common and often ethically acceptable.
@@ 4.1 Framework Overview
Figure 1: Training architecture. Leveraging multi-level contrastive learning loss, we fine-tune a language model (we select XLM-RoBERTa Vamvas and Sennrich (2023), see Appendix E) based on the human, human–LLM and LLM-generated texts, to force the model to reorganize the hidden space, pulling the embeddings within the same author families closer, and pushing the embeddings from different authors farther. We train an encoder that can represent text with distinguishable signals to discern authorship of text.
@@
  • If $x\in P_{2}$ ($x$ is generated by any LLM): Here, with $y_{\textit{LHS}}$ from $P_{3}$ and $y_{\textit{RHS}}$ from $P_{4}$ as defined above, the similarity $\mathrm{S}_{c}(x,y_{\textit{LHS}})$ is generally expected to be equal to $\mathrm{S}_{c}(x,y_{\textit{RHS}})$. Since $x$ can originate from any LLM, it doesn’t inherently possess a stronger connection to the specific LLM family in $P_{3}$ than to the broader human-LLM collaborations represented in $P_{4}$.
@@ 4.2 Multi-level Contrastive Learning
Given a dataset with $N$ samples, each sample is a text unit (paragraph/segment). The $i^{th}$ sample is denoted as $T_{i}$. Per sample, we assigned three-level labels indicating its source:
  • $x_{i}\in\{0,1\}$: if $T_{i}$ is fully LLM-generated, $x_{i}=0$, otherwise $x_{i}=1$;
  • $y_{i}\in\{0,1\}$: if $T_{i}$ is fully human-written, $y_{i}=0$, otherwise $y_{i}=1$;
@@
The encoder $\phi(\cdot)$ represents the sample $T_{i}$ in a $d$-dimensional vector space $\mathbb{R}^{d}$. Then we calculate the cosine similarity between two samples $T_{i}$ and $T_{j}$, denoted by: $\sigma(i,j)=\mathrm{S}_{c}\left(\phi(T_{i}),\phi(T_{j})\right)$.
@@
If $x_{i}=0$, then the text sample is fully LLM-generated. For this case, we do not consider $y_{i}$ since LLM-generated text is considered as non-LLM-collaboration. We can imply that the similarity between two texts written by the same LLM is higher than that of two LLM families. Hence:
@@
For cases where all samples are human–LLM collaborative, two texts created by the same LLM family tend to be more similar than the ones that involve contributions from different families. That is:
@@
Different constraints result in varied sets of positive and negative samples. Based on these sets, contrastive losses are computed at multiple levels. As we declared in Eq (7), each inequality in Eq (6) is denoted as $\mathcal{L}_{q_{i},\varepsilon}$ where $\varepsilon=\overline{1,5}$ respectively. To form Eq (8), we need to add the coefficients $\alpha,\beta,\gamma,\delta$, and $\zeta$ to maintain the balance of multi-level relations.
@@
Therefore, due to the last inequality in Eq (6) only specifying a case ($y_{i,j,k}=1$), and the other cases considering both values for $y$, so we have $\zeta=2\gamma=2\delta$. Also, to maintain the equilibrium, we need to keep $\alpha+\beta=\gamma+\delta+\zeta$. We set $\gamma=\delta=1$, then $\zeta=2,\alpha=\beta=2$. This encourages the model to capture subtle and detailed features from different sources. As a result, it becomes more adept at recognizing variations in writing styles. This capability enhances its accuracy and strengthens its ability to generalize when detecting LLM-generated text.

### v3

@@ Abstract
The growing collaboration between humans and LLMs in generative tasks has introduced new challenges in distinguishing between human-written, LLM-generated, and human–LLM collaborative texts. In this work, we collect a multilingual, multi-domain, multi-generator dataset FAIDSet. We further introduce a fine-grained detection framework, FAID, to classify text into these three categories and to identify the underlying LLM family of the generator. Unlike existing binary classifiers, FAID is built to capture both authorship and model-specific characteristics. Our method combines multi-level contrastive learning with multi-task auxiliary classification to learn subtle stylistic cues. By modeling LLM families as distinct stylistic entities, we adapt to address distributional shifts without retraining on unseen data. Our results demonstrate that FAID outperforms several baselines, particularly improving generalization accuracy across unseen domains and new LLMs, offering a potential solution to improve transparency and accountability in AI-assisted writing. Our data and code are available at https://github.com/mbzuai-nlp/FAID
**footnotetext: Equal contribution.
@@ 1 Introduction
LLMs have evolved from an assistant tool to a creator or initiator, from helping polish papers to initiating proposals and drafting essays, while humans increasingly serve as optimizers and reviewers. In such deeply collaborative human–LLM settings, measuring human contribution becomes challenging, yet clarifying authorship is critical for accountability and transparency, particularly in educational and academic contexts Wang et al. (2025). This work aims to identify human involvement by distinguishing the origin of a given text (in a multilingual context): (a) fully LLM-generated, (b) fully human-written, or (c) collaboratively produced.
@@
Numerous studies have explored multilingual LLM-generated text detection, but most have focused either on binary detection, i.e., human vs. LLM Wang et al. (2024e, d); Su et al. (2023) or on fine-grained detection limited to English Abassy et al. (2024); Koike et al. (2024); Wang et al. (2023); Zhang et al. (2024). Moreover, both struggle with generalization to unseen domains, languages, and LLM generators Wang et al. (2024c); Li et al. (2024).
@@
Our work aims to bridge this gap by (i) collecting a multilingual, multi-domain, and multi-generator dataset, FAIDSet, for fine-grained detection, i.e., identifying a text into three categories: LLM-generated, human-written, and human–LLM collaborative, and (ii) introducing a framework FAID to improve generalization performance.
@@
Our dataset FAIDSet [^1] https://huggingface.co/datasets/ngocminhta/FAIDSet focuses on the academic field, including paper abstracts, student theses, and reports, and contains generation by a variety of families of LLMs, e.g., GPT, Gemini, DeepSeek, and Llama OpenAI (2024); Gemini Team (2024); DeepSeek-AI (2025); Dubey et al. (2024).
@@
Our detection framework, FAID, treats each LLM as a distinct author, learning specific features in the hidden space to differentiate different authors, instead of classifying based on hand-crafted stylistic features. We achieve this by optimizing a language encoder with multi-level author relationship (e.g., the stylistic similarity between texts from the same LLM family is greater than that between a human and an LLM) to capture author-specific distinguishable signals of an input text using contrastive learning, along with the task of fine-tuning a classifier to recognize the input text’s origin. This multitask learning process forces the encoder to reorganize the hidden space so that the representations of texts by the same authors are distributed closer together than those by different authors.
@@
In practice, given that generations produced by LLMs from the same company tend to have similar writing styles due to similar model architecture, training data, and training strategy (see Appendix D for more detail), we consider each LLM family as an “author.” This can also help the detector acquire prior knowledge of future LLMs from the same company. Our experiments show that FAID consistently outperforms other baseline detectors in both in-domain and out-of-domain evaluations. Our contributions can be summarized as follows:
@@
  • We collect a new multilingual, multi-domain, multi-generator dataset for fine-grained LLM-generated text detection with 83,350 examples.
@@
  • We propose a detection framework, FAID, to improve generalization performance in unseen domains and generators by capturing subtle stylistic features in the hidden representation.
@@
  • We show that FAID outperforms baseline detectors, particularly in unseen domains and with unseen generators. Meanwhile, the nature of FAID allows us to assess the stylistic proximity of a given text to other texts in our database, each labeled with ground-truth authorship labels.
@@ 2 Background and Related Work
Advancements in LLMs have fundamentally reshaped the process of text production and refinement. Rather than originating every word independently, people increasingly assume the role of post-editors or reviewers, intervening after an LLM has generated an initial draft. This collaborative paradigm extends across diverse domains, including academic publishing, journalism, education, and social media, thus making hybrid human–AI authorship an emerging norm Cheng et al. (2025); Lee et al. (2022).
Such a shift calls into binary authorship detection: human vs. AI. Texts can be fully human-written, fully AI-generated, or collaborative text (e.g., human-written, AI-polished; AI-generated, human-edited; and deeply mixedArtemova et al. (2025). Addressing these concerns requires fine-grained authorship detection, even tracing it back to a specific LLM family, to assess the extent of human contribution Hutson (2025). Ensuring transparent disclosure of LLM involvement is critical to upholding research integrity and honesty.
In response to these challenges, we collect a new dataset, FAIDSet, and propose a detection framework, FAID, which generalizes well to new domains, languages, and models, achieving consistently high accuracy and reliability.
@@ 2.1 Fine-Grained AI-generated Text Datasets
Many prior studies have explored fine-grained AI-generated text datasets across various forms of human–AI collaboration, e.g., MixSet (Zhang et al., 2024) and Beemo (Artemova et al., 2025). See Appendix A for a discussion of more datasets and detailed information regarding the label space, and the coverage of domains, languages, and LLMs for each one.
However, all these studies were limited to English. A substantial gap remains in the availability of large-scale, fine-grained multilingual LLM-generated text detection datasets. To bridge this gap, we collect a multilingual fine-grained LLM-generated text detection dataset, which encompasses 83k texts generated by the latest LLMs and includes diverse forms of human–LLM collaborative generations. This dataset can facilitate the development of more robust and generalizable detection models that are capable of handling complex multilingual collaborative scenarios.
2.2 Generalization of LLM-Generated Text Detection
A recent study, M4GT-Bench, Wang et al. (2024c) has highlighted a persistent challenge for both binary and fine-grained AI-generated text detection: poor generalization to unseen domains, languages, and generators. Many detection methods have shown a significant drop in performance on out-of-distribution data, underscoring the difficulty of building robust detectors for real-world scenarios and evolving LLM outputs.
Various approaches have been proposed to improve generalization. OUTFOX Koike et al. (2024) leveraged adversarial in-context learning to dynamically generate challenging examples that enhance robustness, but still faced limitations in domain transferability and computational efficiency. LLM-DetectAIve Abassy et al. (2024) adopted fine-grained classification and incorporated domain-adversarial training to reduce overfitting; however, its generalization to unseen domains and generators remains limited, and its current version lacks multilingual support.
SeqXGPT Wang et al. (2023) focused on sentence-level detection by combining log-probabilities with convolutional and self-attention mechanisms, thereby helping capture subtle mixed-content signals and improving generalization across input styles. However, its reliance on specific model features and its limited semantic representation constrained the adaptability to new generators and domains.
Finally, the DeTeCtive framework Guo et al. (2024) introduced multi-level contrastive learning to better capture writing-style diversity and enhance generalizability, especially for out-of-distribution scenarios, but it primarily focused on binary classification and did not fully address human–AI collaborative texts.
@@ 2.3 Contrastive Learning for AI-Generated Text Detection
Contrastive learning has been widely used to improve sentence representations by pulling semantically similar sentences closer together and pushing dissimilar ones apart. SimCSE treats a sentence subjected to dropout noise as a semantically similar counterpart (i.e., a positive pair) and trains the encoder to minimize the distance between the original and the noise sentence. It further leverages natural language inference pairs, considering entailment pairs as positives and contradiction pairs as hard negatives (Gao et al., 2021). It is then trained to maximize the separation between negative pairs in the embedding space. DeCLUTR Giorgi et al. (2021) constructed positive pairs by extracting different spans from the same texts and by sampling negative pairs from different texts.
We adopt the same core idea, but reorganize the latent space by clustering human-written texts by writing style, keeping them distant from LLM-generated texts. Similarly to semantic textual similarity tasks, where sentence similarity ranges from 0 to 5 to reflect varying degrees of semantic overlap, we incorporate ordinal regression into our framework to model the degree of human involvement, ranging from 0 (solely LLM) to 1 (fully human).
Another work, DeTeCtive Guo et al. (2024), also leveraged contrastive learning and was used for binary detection task. Based on a multi-task framework, it was trained to learn style diversity using a multi-level contrastive loss, and an auxiliary task of classifying the source of a given text (human vs. AI) to capture distinguishable signals.
For inference, their pipeline encoded the input text as a hidden vector and used dense retrieval to match the cluster based on stylistic similarity against a database of previously indexed training features. Additionally, instead of retraining the model on new data, they encoded the new data with the trained encoder to obtain embeddings, then added them to the feature database to augment. This largely improved the generalizability to unseen domains and new generators.
However, this approach distinguishes only between two categories of text (human-written vs. LLM-generated) while overlooking the increasingly prevalent class of human–LLM collaborative texts. Our work bridges this gap to enhance generalization performance in fine-grained LLM-generated text detection.
@@ 3 FAIDSet
We collected a new multilingual, multi-domain, and multi-generator LLM-generated text dataset, FAIDSet. It contains texts generated by LLMs, written by humans, and collaborated by both, resulting in a total of 83,350 examples; see Appendix B for more details about FAIDSet.
FAIDSet covers two domains: student theses and paper abstracts, where identifying authorship is critical, across two languages (Vietnamese and English). We collected students’ theses from the database of Hanoi University of Science and Technology, and paper abstracts from arXiv and Vietnam Journals Online [^2] https://vjol.info.vn/ (VJOL).
@@
Diverse Prompt Strategies. We generated data with diverse tones and contexts while ensuring content accuracy. Depending on the data source and context, we crafted prompts to create varied outputs suitable for different real-world scenarios. We generated responses with different tones using prompts such as “You are an IT student…” and “…who are very familiar with abstract writing…”. See Appendix B.2 for the full list.
While FAIDSet does not capture the full natural diversity of in-the-wild LLMs’ outputs, the controlled setup enables us to systematically model stylistic and collaborative signals across multiple languages and families. This makes FAIDSet both a reproducible training resource and a benchmarking corpus: it provides reliable supervision for developing detectors while also serving as a testbed for assessing generalization to more diverse, unseen scenarios.
Quality Control. To avoid bad machine-generated texts, which can introduce remarkably distinguishable signals, we performed quality control by randomly sampling 10–20 instances for each domain, source, and LLM generator. Manual inspection focused on fluency, coherence, and factual plausibility. Overall, the generated texts demonstrated high linguistic quality, with most outputs being fluent and logically reasonable. Nevertheless, occasional issues such as repetitive phrasing, incomplete reasoning, or overuse of formal expressions were observed. In those cases, we adjusted the prompts (e.g., by specifying desired length or style) or refined generation parameters to improve diversity and logical consistency. After these adjustments, the quality across generators and domains was found to be stable and satisfactory.
@@ 4 Methodology
The human–LLM collaborative category involves a range of interactions between humans and LLM systems, such as (a) human-written, LLM-polished, (b) human-initiated, LLM-continued, (c) human-written, LLM-paraphrased, etc. Given the growing variety and complexity (e.g., deeply mixed text) of collaborative patterns, this is not exhaustive. Instead, we consolidate all forms of human–LLM collaboration into a single label to maintain practical simplicity and model generalizability. This reflects real-world usage, where using LLM tools to enhance clarity or expression is increasingly common and often ethically acceptable.
@@ 4.1 Framework Overview
Figure 1: Training architecture. Leveraging multi-level contrastive learning loss, we fine-tune a language model (we select XLM-RoBERTa Vamvas and Sennrich (2023), see Appendix E) based on the human, human–LLM and LLM-generated texts, to force the model to reorganize the hidden space, pulling the embeddings within the same author families closer, and pushing the embeddings from different authors farther. We train an encoder that can represent text with distinguishable signals to discern the authorship of text.
@@
  • If $x\in P_{2}$ ($x$ is generated by any LLM): Here, with $y_{\textit{LHS}}$ from $P_{3}$ and $y_{\textit{RHS}}$ from $P_{4}$ as defined above, the similarity $\mathrm{S}_{c}(x,y_{\textit{LHS}})$ is generally expected to be equal to $\mathrm{S}_{c}(x,y_{\textit{RHS}})$. Since $x$ can originate from any LLM, it does not inherently possess a stronger connection to the specific LLM family in $P_{3}$ than to the broader human-LLM collaborations represented in $P_{4}$.
@@ 4.2 Multi-level Contrastive Learning
Given a dataset with $N$ examples, each example is a text unit (paragraph/segment). The $i^{th}$ record is denoted as $T_{i}$. Per record, we assigned three-level labels indicating its source:
  • $x_{i}\in\{0,1\}$: if $T_{i}$ is a fully LLM-generated text, $x_{i}=0$, otherwise $x_{i}=1$;
  • $y_{i}\in\{0,1\}$: if $T_{i}$ is a fully human-written text, $y_{i}=0$, otherwise $y_{i}=1$;
@@
The encoder $\phi(\cdot)$ represents the text $T_{i}$ in a $d$-dimensional vector space $\mathbb{R}^{d}$. Then we calculate the cosine similarity between two texts $T_{i}$ and $T_{j}$, denoted by: $\sigma(i,j)=\mathrm{S}_{c}\left(\phi(T_{i}),\phi(T_{j})\right)$.
@@
If $x_{i}=0$, then the text is fully LLM-generated. For this case, we do not consider $y_{i}$ since LLM-generated text is considered as non-LLM-collaboration. We can imply that the similarity between two texts written by the same LLM is higher than that of two LLM families. Hence:
@@
For cases where all samples are human–LLM collaborative, two texts created by the same LLM family tend to be more similar than such that involve contributions from different LLM families. That is:
@@
Different constraints yield different sets of positive and negative samples. Based on these sets, contrastive losses are computed at multiple levels. As we declared in Eq (7), each inequality in Eq (6) is denoted as $\mathcal{L}_{q_{i},\varepsilon}$ where $\varepsilon=\overline{1,5}$ respectively. To form Eq (8), we need to add the coefficients $\alpha,\beta,\gamma,\delta$, and $\zeta$ to maintain the balance of multi-level relations.
@@
Due to the last inequality in Eq (6) only specifying a case ($y_{i,j,k}=1$), and the other cases considering both values for $y$, we have $\zeta=2\gamma=2\delta$. Also, to maintain the equilibrium, we need to keep $\alpha+\beta=\gamma+\delta+\zeta$. We set $\gamma=\delta=1$, then $\zeta=2,\alpha=\beta=2$. This encourages the model to capture subtle and detailed features from different sources. As a result, it becomes more adept at recognizing variations in writing styles. This capability enhances accuracy and strengthens generalizability when detecting LLM-generated text.